# Synergistic Dynamical Decoupling and Circuit Design for Enhanced Algorithm Performance on Near-Term Quantum Devices

### v1

@@ III.2.2 DD sequences
Figure 9: Comparison of CPMG and XY4 sequences across seven IBM QPUs ibmq_mumbai, ibmq_kolkata, ibm_cairo, ibmq_ehningen, ibm_kyoto, ibm_cusco, and ibm_brisbane. Each line represents a linear fit of the data.
@@
$\mathrm{NAR}_{\mathrm{B}}^{\mathrm{CPMG}}$ 0.610 $y=-0.050x+0.986$ $-0.749$ 0
$\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{CPMG}}$ 0.699 $y=-0.044x+1.029$ $-0.802$ 0
$\mathrm{NAR}_{\mathrm{B}}^{\mathrm{XY4}}$ 0.6 $y=-0.052x+0.992$ $-0.752$ 0
$\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{XY4}}$ 0.697 $y=-0.043x+1.019$ $-0.750$ 0
$\mathrm{NSP}_{\mathrm{B}}^{\mathrm{CPMG}}$ 0.525 $y=-0.061x+0.986$ $-0.613$ 0
$\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{CPMG}}$ 0.629 $y=-0.051x+1.009$ $-0.513$ 0.00001
$\mathrm{NSP}_{\mathrm{B}}^{\mathrm{XY4}}$ 0.509 $y=-0.057x+0.940$ $-0.591$ 0
$\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{XY4}}$ 0.648 $y=-0.044x+0.975$ $-0.373$ 0.00148
@@
$\ln(\tau)$ (CPMG) 10.589 $y=\phantom{-}0.165x+9.354$ $8510$
$\ln(\tau)$ (XY4) 10.586 $y=\phantom{-}0.163x+9.360$ $8580$
This section evaluates the performance of two DD sequences, CPMG and XY4, for mitigating decoherence errors during QAOA execution with CX implementation. The evaluation leverages data from seven IBM QPUs. XY4, which employs four single qubit pulses, might be more effective for qubits with extended idle times than CPMG, which utilizes two single qubit pulses.
Figures 9(a) and 9(b) present the results for NAR and NSP, respectively. The data reveal that both CPMG and XY4 contribute to improved algorithm performance. While CPMG and XY4 achieve comparable NAR values, XY4 exhibits better performance in terms of NSP for a larger number of qubits. However, the results for XY4 also display significant fluctuations. $\Delta_{\mathrm{NAR}}$ and $\Delta_{\mathrm{NSP}}$, as shown in Figs. 9(c) and 9(d), demonstrate comparable performance between CPMG and XY4 for a small number of qubits, while XY4 appears to be slightly more effective for a larger number of qubits. A key observation lies in the EMSR values. The reported $\mathrm{EMSR}_{\mathrm{AR}}$ and $\mathrm{EMSR}_{\mathrm{SP}}$ for CPMG are 84.29% and 75.71%, respectively, compared to 67.14% and 64.29% for XY4. This suggests a higher degree of robustness for the CPMG sequence. Figures 9(e) and 9(f) illustrate the circuit fidelity and schedule duration, respectively, with comparable results observed for both CPMG and XY4.
As before, the data are fitted with a linear function. The resulting parameters for CPMG and XY4 sequences are summarized in Table 6. The coefficients of the linear function suggest that both CPMG and XY4 effectively suppress the decrease of NAR and NSP as the qubit count increases. The correlation coefficients and $p$-values suggest a stronger correlation between $\Delta_{\mathrm{NAR}}^{\mathrm{XY4}}$ and qubit count compared to $\Delta_{\mathrm{NAR}}^{\mathrm{CPMG}}$, $\Delta_{\mathrm{NSP}}^{\mathrm{XY4}}$, and $\Delta_{\mathrm{NSP}}^{\mathrm{CPMG}}$.
The results indicate that in principle, DD is generally recommended. The utilization of DD sequences expands the range of circuits where acceptable results can be achieved. For instance, QAOA can achieve an NAR value greater than 0.6 for a larger number of qubits when DD is employed. However, the effectiveness and robustness of DD sequences can vary. While XY4 yields slightly greater enhancements in algorithm performance, as indicated by NAR and NSP, CPMG demonstrates higher robustness, as evidenced by its higher EMSR values. This highlights the importance of considering not only performance improvements but also the trade-off with mitigation robustness when selecting a DD sequence.
@@ III.2.3 Optimization levels
Figure 10: Comparison of optimization level 1 (Opt1) and optimization level 3 (Opt3) across five IBM QPUs ibmq_kolkata, ibm_cairo, ibmq_ehningen, ibm_cusco, and ibm_kyoto. The CPMG sequence is used for all data points. Each line represents a linear fit of the data.
We investigate the influence of optimization levels on the performance and DD effectiveness using five IBM QPUs. Two optimization levels are considered: level 1 (Opt1) and level 3 (Opt3), representing the default and highest settings, respectively. The Qiskit’s transpiler offers four optimization levels, ranging from 0 to 3. Level 0 performs no optimization, while level 3 performs the most extensive optimization. The default setting in Qiskit’s transpiler is level 1, which performs light optimization. In comparison, level 2 performs heavy optimization. It is worth noting that different optimization levels do not affect the number of two qubit gates in our demonstration. This is because we are considering benchmark circuits that have already undergone the AOQMAP approach [44], which effectively ensures adherence to connectivity constraints and eliminates the need for additional SWAP gates. In our case, different optimization levels influence the selected qubits for circuit execution and the count of single qubit gates, which affects the circuit fidelity and schedule duration. The CPMG sequence and CX implementation of QAOA are employed throughout this analysis.
@@
$\mathrm{NAR}_{\mathrm{B}}^{\mathrm{Opt1}}$ 0.554 $y=-0.052x+0.946$ $-0.891$ 0
$\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{Opt1}}$ 0.657 $y=-0.046x+1.004$ $-0.886$ 0
$\mathrm{NAR}_{\mathrm{B}}^{\mathrm{Opt3}}$ 0.509 $y=-0.050x+0.885$ $-0.596$ 0
$\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{Opt3}}$ 0.589 $y=-0.055x+0.998$ $-0.667$ 0
$\mathrm{NSP}_{\mathrm{B}}^{\mathrm{Opt1}}$ 0.525 $y=-0.061x+0.986$ $-0.613$ 0
$\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{Opt1}}$ 0.629 $y=-0.051x+1.009$ $-0.513$ 0.00001
$\mathrm{NSP}_{\mathrm{B}}^{\mathrm{Opt3}}$ 0.536 $y=-0.060x+0.988$ $-0.586$ 0.00001
$\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{Opt3}}$ 0.606 $y=-0.050x+0.981$ $-0.471$ 0.00056
@@
$\Delta_{\mathrm{NSP}}^{\mathrm{Opt1}}$ 0.08 $y=-0.004x+0.113$ $-0.065$ 0.65396
@@
$\ln(\tau)$ (Opt1) 10.540 $y=\phantom{-}0.183x+9.168$ $9190$
$\ln(\tau)$ (Opt3) 10.475 $y=\phantom{-}0.172x+9.186$ $9010$
Figure 10(a) and 10(b) demonstrate that Opt3 with DD sequences achieves the highest overall performance in terms of NAR and NSP. This is followed by Opt1 with DD sequences ($\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{Opt1}}$ and $\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{Opt1}}$), highlighting the importance of DD sequences in enhancing performance for both optimization levels. Without DD sequences, Opt3 outperforms Opt1 for all tested qubit counts in terms of NAR. However, for NSP, Opt3 exhibits an advantage only for a small number of qubits, with comparable performance achieved at larger qubit counts. The effectiveness of DD sequences is further explored in Figs. 10(c) and 10(d). Figure 10(c) shows that the average improvement in NAR due to DD sequences ($\Delta_{\mathrm{NAR}}$) is generally higher for Opt1 compared to Opt3. This suggests a stronger mitigating effect of DD sequences on performance degradation for Opt1. Similarly, Fig. 10(d) reveals that DD sequences are initially more beneficial for Opt1 in terms of NSP, with the advantage shifting towards Opt3 for larger qubit counts. Furthermore, the reported average $\mathrm{EMSR}_{\mathrm{AR}}$ and $\mathrm{EMSR}_{\mathrm{SP}}$ are 92% and 74% for Opt1, respectively, compared to 80% and 60% for Opt3. This indicates a potential advantage of DD sequences in terms of robustness for Opt1. It is important to note that Opt3 exhibits higher circuit fidelity (Fig. 10(e)) and shorter schedule duration (Fig. 10(f)), which are crucial for high algorithm performance.
The detailed linear fit parameters for the optimization levels are provided in Table 7. When DD sequences are applied, the decay of NAR is suppressed for Opt1, and the decay of NSP is suppressed for both Opt1 and Opt3 as the number of qubits increases. However, for Opt3, DD sequences appear to worsen the decay of NAR. This trend is also reflected in the corresponding correlation coefficients. These observations indicate that the effectiveness of DD sequences may decrease as the number of qubits increases, possibly due to errors accumulating within the sequences themselves. The correlation coefficient $C_{r}$ is very low and the $p$-value is very high for $\Delta_{\mathrm{NSP}}^{\mathrm{Opt1}}$, suggesting a very weak or no correlation between NSP with Opt1 and the number of qubits. Furthermore, Opt1 shows a stronger correlation with schedule duration, while Opt3 is more strongly correlated with circuit fidelity.
@@ IV Discussion and Conclusions
Our comprehensive study, conducted on eight IBM quantum devices, reveals that the application of dynamical decoupling (DD) sequences can significantly enhance the performance and robustness of algorithms on near term quantum devices. However, the effectiveness of DD sequences varies depending on hardware and algorithmic factors. A key finding is the observed inverse relationship between DD effectiveness and the original performance of algorithms without error mitigation. This implies that algorithms with higher inherent performance (measured without DD sequences) exhibit lower DD effectiveness. For instance, ECR-based QPUs offer superior native performance but reduced DD effectiveness compared to CX-based QPUs. Similarly, CZ implementation of QAOA exhibits higher native algorithm performance but lower average DD effectiveness compared to CX implementation. Moreover, optimization level 3 produces higher algorithm performance, but level 1 exhibits higher DD effectiveness. This inverse behavior can be attributed to the fact that algorithms with high performance typically have a lower intrinsic error rate, including the decoherence errors targeted by DD sequences. Furthermore, the introduction of DD pulse sequences can lead to new gate operation errors that decrease algorithm performance and potentially limit their effectiveness for certain algorithms.
@@
DD sequences are typically more effective for algorithms with lower fidelity and longer schedule duration, but their impact on approximation ratio and success probability differs. The results indicate that while algorithms with lower circuit fidelity struggle to achieve high approximation ratio values, the application of DD sequences allows for achieving the simulated success probability. This finding suggests a potential methodology for obtaining the desired probabilities by studying different algorithm implementations and DD sequences, and selecting the measure that minimizes system energy.
@@
One significant reason for the higher performance of algorithms on ECR-based QPUs is the inherently higher circuit fidelity and shorter schedule duration. However, another potential factor could be the inherent advantages of ECR gates. In these devices, ECR gates are only allowed for one direction, meaning that any two-qubit gate is directly decomposed into ECR gates with the same direction on the qubit pair. While the hardware-native CX gate also has a native direction, the reverse direction is supported, requiring additional single qubit gates and the hardware native CX gate for implementation. Directly decomposing algorithms into one-directional ECR gates could be more advantageous than using CX gates with bidirectional capability at the gate level and then transforming them into native CX gates at the hardware pulse level. A similar trend is observed for CZ implementation, which produces a higher algorithm performance than CX implementation. The CX gate is directed, whereas the CZ gate is undirected, allowing for the decomposition of all two qubit gates with one directed CZ gate on one or more qubit pairs. Utilizing gates in the same direction could lead to a higher symmetry in the circuit, both in terms of single and two qubit gates, potentially contributing to error suppression and improved algorithmic performance. This highlights the importance of considering native gates at the pulse level and maintaining circuit structure symmetry during algorithm design.
@@
In conclusion, this study demonstrates the significant impact of several factors on algorithm performance and the effectiveness of error mitigation across eight IBM QPUs. These factors include circuit fidelity, schedule duration, choice of hardware native gates, algorithm implementations, types of DD sequences, and optimization levels. Despite minimal performance variations between Carr-Purcell-Meiboom-Gill (CPMG) and XY4 sequences, XY4 exhibits a slight advantage in success probability for larger qubit counts. However, CPMG achieves a higher overall error mitigation success rate, suggesting potentially greater robustness. While the results highlight the general enhancement of algorithm performance and robustness through the use of DD sequences, achieving significantly improved performance relies more critically on factors such as high quality native gates of QPUs, symmetric algorithm implementation, and effective circuit optimization techniques. Therefore, a holistic approach that considers both hardware characteristics and software optimization strategies is important when designing quantum algorithms to maximize reliability and efficacy. This study underscores the importance of hardware considerations and circuit design in enhancing algorithm performance and robustness using DD sequences. These insights guide the development and optimization of other quantum applications. Future research directions include investigating the interplay between these factors when combining DD sequences with other error mitigation strategies, such as zero-noise extrapolation. Additionally, exploring the relationship between circuit symmetry and its effectiveness in suppressing errors could provide valuable insights. Extending this analysis to other quantum algorithms, such as the variational quantum eigensolver (VQE) and protocols for preparing Greenberger-Horne-Zeilinger (GHZ) states, holds promise for revealing the broader applicability of these findings.

### v2

@@ III.2.2 DD Sequences
This section evaluates the performance of two DD sequences, CPMG and XY4, for mitigating decoherence errors during QAOA execution with CX implementation. The evaluation leverages data from seven IBM QPUs. XY4, which employs four single-qubit pulses, might be more effective for qubits with extended idle times than CPMG, which utilizes two single-qubit pulses. While DD sequences may introduce crosstalk errors, the use of the same transpiled circuits minimizes the impact of this potential crosstalk on our evaluation.
@@
Figure 9: Comparison of CPMG and XY4 sequences across seven IBM QPUs ibmq_mumbai, ibmq_kolkata, ibm_cairo, ibmq_ehningen, ibm_kyoto, ibm_cusco, and ibm_brisbane: (a) NAR, (b) NSP, (c) $\Delta_{\mathrm{\scriptscriptstyle NAR}}$, (d) $\Delta_{\mathrm{\scriptscriptstyle NSP}}$, (e) circuit fidelity, and (f) $\ln(\tau/\mathrm{dt})$. Each line represents a linear fit of the data.
@@
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle CPMG}}$ 0.610 $y=-0.050x+0.986$ $-0.749$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CPMG}}$ 0.699 $y=-0.044x+1.029$ $-0.802$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle XY4}}$ 0.6 $y=-0.052x+0.992$ $-0.752$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle XY4}}$ 0.697 $y=-0.043x+1.019$ $-0.750$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle CPMG}}$ 0.525 $y=-0.061x+0.986$ $-0.613$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CPMG}}$ 0.629 $y=-0.051x+1.009$ $-0.513$ 0.00001
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle XY4}}$ 0.509 $y=-0.057x+0.940$ $-0.591$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle XY4}}$ 0.648 $y=-0.044x+0.975$ $-0.373$ 0.00148
@@
$\ln(\tau/\mathrm{dt})$ (CPMG) 10.589 $y=\phantom{-}0.165x+9.354$ $8510$
$\ln(\tau/\mathrm{dt})$ (XY4) 10.586 $y=\phantom{-}0.163x+9.360$ $8580$
Figures 9(a) and (b) show that both CPMG and XY4 contribute to improved algorithm performance. While CPMG and XY4 achieve comparable NAR values, XY4 exhibits a better NSP for a larger number of qubits, but with noticeable fluctuations. CPMG and XY4 demonstrate comparable DD effectiveness for a small number of qubits (Figs. 9(c) and (d)), with XY4 showing a slight advantage for larger qubit counts. However, CPMG exhibits greater robustness, as evidenced by its higher $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle AR}}$ (84.29%) and $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle SP}}$ (75.71%) compared to XY4’s values (67.14% and 64.29%, respectively). Figures 9(e) and (f) illustrate comparable circuit fidelity and schedule duration for both CPMG and XY4. As before, the data are fitted with a linear function. The resulting parameters are summarized in Table 6. The coefficients of the linear function suggest that both CPMG and XY4 effectively suppress the decrease in NAR and NSP as the qubit count increases. The correlation coefficients and $p$-values suggest a stronger correlation between $\Delta_{\mathrm{\scriptscriptstyle NAR}}^{\mathrm{\scriptscriptstyle XY4}}$ and qubit count compared to $\Delta_{\mathrm{\scriptscriptstyle NAR}}^{\mathrm{\scriptscriptstyle CPMG}}$, $\Delta_{\mathrm{\scriptscriptstyle NSP}}^{\mathrm{\scriptscriptstyle XY4}}$, and $\Delta_{\mathrm{\scriptscriptstyle NSP}}^{\mathrm{\scriptscriptstyle CPMG}}$.
The results indicate that DD sequences are generally recommended for improving circuit performance. They allow for achieving acceptable results in a wider range of circuits. For instance, the QAOA with DD can reach higher NAR values for more qubits. However, the effectiveness and robustness of DD sequences can vary. While XY4 offers slightly better performance improvements in terms of NAR and NSP, CPMG demonstrates higher robustness as measured by EMSR. This highlights the importance of considering both performance gains and mitigation robustness when choosing a DD sequence.
@@ III.2.3 Optimization Levels
We investigate the influence of optimization levels on the performance and DD effectiveness using five IBM QPUs. Two optimization levels within Qiskit’s transpiler are considered: level 1 (Opt1) and level 3 (Opt3), representing the default and highest settings, respectively. It is worth noting that different optimization levels do not affect the number of two-qubit gates in our demonstration. This is because we are considering benchmark circuits that have already undergone the AOQMAP approach [42], which effectively ensures adherence to connectivity constraints and eliminates the need for additional SWAP gates. In our case, different optimization levels influence the selected qubits for circuit execution and the count of single-qubit gates, which affects the circuit fidelity and schedule duration. The CPMG sequence and CX implementation of the QAOA are employed throughout this analysis.
@@
Figure 10: Comparison of optimization level 1 (Opt1) and optimization level 3 (Opt3) across five IBM QPUs ibmq_kolkata, ibm_cairo, ibmq_ehningen, ibm_cusco, and ibm_kyoto: (a) NAR, (b) NSP, (c) $\Delta_{\mathrm{\scriptscriptstyle NAR}}$, (d) $\Delta_{\mathrm{\scriptscriptstyle NSP}}$, (e) circuit fidelity, and (f) $\ln(\tau/\mathrm{dt})$. The CPMG sequence is used for all data points. Each line represents a linear fit of the data.
@@
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle Opt1}}$ 0.554 $y=-0.052x+0.946$ $-0.891$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle Opt1}}$ 0.657 $y=-0.046x+1.004$ $-0.886$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle Opt3}}$ 0.618 $y=-0.048+0.982$ $-0.720$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle Opt3}}$ 0.688 $y=-0.045x+1.023$ $-0.770$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle Opt1}}$ 0.509 $y=-0.050x+0.885$ $-0.596$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle Opt1}}$ 0.589 $y=-0.055x+0.998$ $-0.667$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle Opt3}}$ 0.536 $y=-0.060x+0.988$ $-0.586$ 0.00001
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle Opt3}}$ 0.606 $y=-0.050x+0.981$ $-0.471$ 0.00056
@@
$\Delta_{\mathrm{\scriptscriptstyle NSP}}^{\mathrm{\scriptscriptstyle Opt1}}$ 0.08 $y=-0.004x+0.113$ $-0.065$ 0.65396
@@
$\ln(\tau/\mathrm{dt})$ (Opt1) 10.540 $y=\phantom{-}0.183x+9.168$ $9190$
$\ln(\tau/\mathrm{dt})$ (Opt3) 10.475 $y=\phantom{-}0.172x+9.186$ $9010$
Figures 10(a) and (b) demonstrate that Opt3 with DD sequences achieves the highest overall performance in terms of NAR and NSP, followed by Opt1 with DD sequences. Without error mitigation, Opt3 outperforms Opt1 for all tested qubit counts in terms of NAR, whereas, for NSP, Opt3 exhibits an advantage only for a small number of qubits, with comparable performance achieved at larger qubit counts. The average improvement in NAR due to DD ($\Delta_{\mathrm{\scriptscriptstyle NAR}}$) is generally higher for Opt1 compared to Opt3 (Fig. 10(c)). However, for NSP, DD initially benefits Opt1 more, but this advantage shifts toward Opt3 for larger qubit counts (Fig. 10(d)). Furthermore, the reported average $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle AR}}$ and $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle SP}}$ are 92% and 74% for Opt1, respectively, compared to 80% and 60% for Opt3, suggesting a higher robustness of DD for Opt1. It is important to note that Opt3 exhibits higher circuit fidelity (Fig. 10(e)) and a shorter schedule duration (Fig. 10(f)), which are crucial for high algorithm performance.
The detailed linear fit parameters for the optimization levels are provided in Table 7. While DD effectively mitigates the decrease in NAR for both Opt1 and Opt3 with increasing qubits, its impact on NSP differs. For Opt3, DD suppresses the decrease in NSP, whereas, for Opt1, it appears to exacerbate the decay. The correlation between NSP and qubit count for Opt1 is very weak (low $C_{r}$ and high $p$-value), suggesting minimal influence from qubit count on NSP for this optimization level. In contrast, Opt1 prioritizes schedule duration, exhibiting a stronger correlation with qubit count, while Opt3 prioritizes circuit fidelity, showing a stronger correlation between fidelity and qubit count.
@@ IV Discussion and Conclusions
Our comprehensive study, conducted on eight IBM quantum devices, reveals that the application of dynamical decoupling (DD) sequences can significantly enhance the performance and robustness of algorithms on near-term quantum devices. However, the effectiveness of DD sequences varies depending on hardware and algorithmic factors. A key finding is the observed inverse relationship between DD effectiveness and the original performance of algorithms without error mitigation. This implies that algorithms with higher inherent performance (measured without DD sequences) exhibit lower DD effectiveness. For instance, ECR-based QPUs offer superior native performance but reduced DD effectiveness compared to CX-based QPUs. Similarly, the CZ implementation of a QAOA exhibits higher native algorithm performance but lower average DD effectiveness compared to CX implementation. Moreover, optimization level 3 produces higher algorithm performance, but level 1 exhibits higher DD effectiveness. This inverse behavior can be attributed to the fact that algorithms with high performance typically have a lower intrinsic error rate, including the decoherence errors targeted by DD sequences. Furthermore, the introduction of DD pulse sequences can lead to new gate operation errors that decrease algorithm performance and potentially limit their effectiveness for certain algorithms.
@@
DD sequences are typically more effective for algorithms with lower fidelity and a longer schedule duration, but their impact on approximation ratio and success probability differs. The results indicate that while algorithms with lower circuit fidelity struggle to achieve high approximation ratio values, the application of DD sequences allows for achieving the simulated success probability. This finding suggests a potential methodology for obtaining the desired probabilities by studying different algorithm implementations and DD sequences and selecting the measure that minimizes system energy.
[…]
Table 8: Impact of hardware and algorithm factors on DD effectiveness and robustness.
Factor Mean Slope Coefficient EMSR
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}$ $\Delta_{\mathrm{\scriptscriptstyle NAR}}$ $\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}$ $\Delta_{\mathrm{\scriptscriptstyle NSP}}$ $\phantom{-}\Delta_{\mathrm{\scriptscriptstyle NAR}}$ $\phantom{-}\Delta_{\mathrm{\scriptscriptstyle NSP}}$ $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle AR}}$ $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle SP}}$
Circuit fidelity 0.610 0.077 0.556 0.075 $-0.065$ $-0.021$ 85.55% 66.80%
[…]
@@
One significant reason for the higher performance of algorithms on ECR-based QPUs is the inherently higher circuit fidelity and shorter schedule duration. However, another potential factor could be the inherent advantages of ECR gates. In these devices, ECR gates are only allowed for one direction, meaning that any two-qubit gate is directly decomposed into ECR gates with the same direction on the qubit pair. While the hardware-native CX gate also has a native direction, the reverse direction is supported, requiring additional single-qubit gates and the hardware-native CX gate for implementation. Directly decomposing algorithms into one-directional ECR gates could be more advantageous than using CX gates with bidirectional capability at the gate level and then transforming them into native CX gates at the hardware pulse level. A similar trend is observed for CZ implementation, which produces a higher algorithm performance than CX implementation. The CX gate is directed, whereas the CZ gate is undirected, allowing for the decomposition of all two-qubit gates with one directed CZ gate on one or more qubit pairs. Utilizing gates in the same direction could lead to a higher symmetry in the circuit, both in terms of single- and two-qubit gates, potentially contributing to error suppression and improved algorithmic performance. This highlights the importance of considering native gate properties at the pulse level and maintaining circuit structure symmetry during algorithm design. Our findings hold broad applicability across various quantum algorithms, including the variational quantum eigensolver (VQE) [52, 53]. Prioritizing native gates and maintaining circuit symmetry during VQE and other quantum algorithm executions can enhance performance and mitigate errors across diverse quantum computing platforms.
@@
In conclusion, this study demonstrates the significant impact of several factors on algorithm performance and the effectiveness of error mitigation across eight IBM QPUs. These factors include circuit fidelity, schedule duration, choice of hardware-native gates, algorithm implementations, types of DD sequences, and optimization levels. Despite minimal performance variations between Carr–Purcell–Meiboom–Gill (CPMG) and XY4 sequences, XY4 exhibits a slight advantage in success probability for larger qubit counts. However, CPMG achieves a higher overall error mitigation success rate, suggesting potentially greater robustness. While the results highlight the general enhancement of algorithm performance and robustness through the use of DD sequences, achieving significantly improved performance relies more critically on factors such as high-quality native gates of QPUs, symmetric algorithm implementation, and effective circuit optimization techniques. Therefore, a holistic approach that considers both hardware characteristics and software optimization strategies is important when designing quantum algorithms to maximize reliability and efficacy. This study underscores the importance of hardware considerations and circuit design in enhancing algorithm performance and robustness using DD sequences. These insights guide the development and optimization of other quantum applications. Future research directions include investigating the interplay between these factors when combining DD sequences with other error mitigation strategies, such as zero-noise extrapolation [54, 55]. Additionally, exploring the relationship between circuit symmetry and its effectiveness in suppressing errors could provide valuable insights. Extending this analysis to other quantum algorithms, such as the VQE and protocols for preparing Greenberger–Horne–Zeilinger (GHZ) states, holds promise for revealing the broader applicability of these findings.